OpenStudy (swaqqout_kid1):

Solve for x. x = -35 x = -28 x = -21 x = 44 @ttb123456789

OpenStudy (anonymous):

did you medaled me in the previous one?

OpenStudy (swaqqout_kid1):



OpenStudy (swaqqout_kid1):

yea

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

is it x/4+9=2?

OpenStudy (swaqqout_kid1):

yea

OpenStudy (anonymous):

ok then x/4=2-9 x=(2-9)*4

OpenStudy (swaqqout_kid1):

-28

OpenStudy (anonymous):

-7*4=-28 YES